OpenStudy (anonymous):

Perform the division. Write the answer with positive exponents. 12x^10+30x^7 over 3x^4

OpenStudy (accessdenied):

\[ \frac{12x^{10} + 30x^7}{3x^4} = \frac{12x^{10}}{3x^4} + \frac{30x^7}{3x^4} \] We can subtract the variables' exponents, and simplify 12/3 and 30/3. 12/3= 4, 30/3 = 10 \[ 4x^{10 - 4} + 10x^{7-4} = 4x^{6} + 10x^{3} \]

OpenStudy (anonymous):

\[4m ^{2}+9m-28 \over m+4 \] Would that be done the same way??

OpenStudy (accessdenied):

since that one doesn't have a monomial in the denominator, we'd have to use polynomial long division or synthetic division (since its a linear expression with leading coefficient 1)

OpenStudy (anonymous):

So how would we do it with numbers? The words go over my head, sorry!!

OpenStudy (accessdenied):

Synthetic Division goes like this: